How many gallons of cream that is 23% butterfat and milk that is 3% butterfat must a dairy farmer mix to make a 30 gallon mixture of milk that is 4% butterfat?
Answer by stanbon(75887)

How many gallons of cream that is 23% butterfat and milk that is 3% butterfat must a dairy farmer mix to make a 30 gallon mixture of milk that is 4% butterfat?
--------------------
Equation:
fat + fat = fat
0.23x + 0.03(30-x) = 0.04*30
----
Multiply thru by 100 to get:
23x + 3*30-3x = 4*30
20x = 30
x = 3/2 gallons (amt. of 23% cream needed)
=================
Cheers,
Stan H.
=================